A trolley of mass 7 kg is moving with the speed of 1.5 m/s. Calculate the kinetic energy of the trolley.


1
Expert's answer
2021-08-29T16:57:34-0400
"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot7\ kg\cdot(1.5\ \dfrac{m}{s})^2=7.87\ J."